OpenStudy (anonymous):

At 1 p.m., Lin records an outside temperature of –14.8°F. At 3 p.m., she records the outside temperature and notices the temperature had increased by 22.6 degrees. What was the temperature that she recorded at 3 p.m.? _______________________degrees F I think it is either 7.8 or 37.4, i need help, which one is it?

Vocaloid (vocaloid):

"increased" means add -14.8 + 22.6 = ?

OpenStudy (anonymous):

7.8

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

anyone else agree

OpenStudy (anonymous):

i have a feeling it is 37.8

OpenStudy (anonymous):

i mean 37.4, my bad

OpenStudy (anonymous):

wait, it is 7.8 ok

Vocaloid (vocaloid):

are you still confused? would you like me to explain?

OpenStudy (anonymous):

yes plz @Vocaloid

Vocaloid (vocaloid):

well, our starting temperature is -14.8, correct? the problem says that the temperature increases by 22.6 degrees, which means we need to add 22.6 to -14.8 so -14.8 + 22.6 = 7.8 degrees

OpenStudy (mathstudent55):

You need to add: \(-14.8 + 22.6\) You are adding a positive number and a negative number. This is how you do it. First, take the absolute value of both numbers. This means, make both numbers positive, and ignore the negative sign of one of them. Then subtract the smaller number form the larger number. This is the way you always subtracted two numbers before you ever heard of negative numbers. Then to decide whether the final answer is positive or negative, look at the original numbers. If the larger original number (again ignoring signs) is positive, then the answer is positive. If the larger original number (ignoring the sign) is negative, the answer is negative.

OpenStudy (anonymous):

ok, i see, thx @Vocaloid , and @mathstudent55

OpenStudy (anonymous):

so 7.8, thx a bunch u 2

OpenStudy (mathstudent55):

Let's do your problem using the rule above. Your numbers are -14.8 (a negative number) 22.6 (a positive number) We are adding a positive number and a negative number. The first step is ignore the negative sign and think of only 14.8 and 22.6 Now subtract the smaller number form the larger number: 22.6 - 14.8 = 7.8 We know the number in the answer is 7.8. Now we need to find out if the answer will be 7.8 or -7.8. To decide, look again at the original numbers. -14.8 and 22.6 If you take the negative sign off -14.8, you have 14.8. Compare 14.8 with 22.6. Since 22.6 is larger than 14.8, the final answer will have the same sign as 22.6 had originally. Originally, 22.6 was positive, so the answer is positive. -14.8 + 22.6 = 7.8

OpenStudy (anonymous):

wow thats a mouthfull thx